OpenStudy (anonymous):

State the missing steps and reasons to this solution of 3(x+4)=18 a) 3(x+4)=18 b) Distributive property c) 3x+12-12=18-12 d) 3x+0=18-12 e) 3x=18-12 f) Simplify g) 3x/3= 6/3 h) 1x=6/3 i) x=6/3 j) x=2

jimthompson5910 (jim_thompson5910):

hint: the distributive property is a*(b+c) = a*b + a*c

OpenStudy (anonymous):

For b it would be 3x+4=18.

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

c?

OpenStudy (utterly_confuzzled):

B would be 3*x + 3*4 =18 but simplified

jimthompson5910 (jim_thompson5910):

yes or 3x + 12 = 18

OpenStudy (anonymous):

is this 4 a test @AppleBVB

OpenStudy (anonymous):

No, this is not for a test.

OpenStudy (anonymous):

For c it would be to subtract 12 from both sides. For g it would be to divide by 3. I dont know what to do for d, e, h, i, and j.

OpenStudy (anonymous):

No, this is for a test; this is just a warm up problem from school.

OpenStudy (utterly_confuzzled):

D is just showing that once you subtract 12, there's 0 left to add to 3x E is just taking away 0 as a place holder H they are just showing that 3x/3 is equal to 1x I is taking away one since it's unnecessary to write it J is the final answer when 6 is divided by 3 Is that what you were needing??